OpenStudy (anonymous):

Write an expression for the general or nth term, an, for the geometric sequence. 1, -4, 16, -64, . . . A. an = 1 • (-4)n-1 B. an = 1 • (-4)n C. an = 1 • 4n-1 D. an = -1 • 4n

hartnn (hartnn):

do you know what is common ratio in geometric sequence ??

OpenStudy (anonymous):

not really

hartnn (hartnn):

common ratio (r) is the ratio of next term to current term in a geometric sequence, like r = 2nd term/ 1st term = -4/1 = -4 here got this ?

hartnn (hartnn):

no ? any doubts ?

OpenStudy (anonymous):

YA WHAT IS THE ANSWER

hartnn (hartnn):

why you just need the answer ? don't want to learn how to get it ?

OpenStudy (anonymous):

I STILL DONT UNDERSTAND BUT WHAT EVER

hartnn (hartnn):

is -4/1 = 16/-4 = -64/16 ?